# Error Broadcast and Decorrelation as a Potential Artificial and Natural Learning Mechanism

### v2

@@ Abstract
We introduce Error Broadcast and Decorrelation (EBD), a novel learning framework for neural networks that addresses credit assignment by directly broadcasting output errors to individual layers, circumventing weight transport of backpropagation. EBD is rigorously grounded in the stochastic orthogonality property of Minimum Mean Square Error estimators. This fundamental principle states that the error of an optimal estimator is orthogonal to functions of the input. Guided by this insight, EBD defines layerwise loss functions that directly penalize correlations between layer activations and output errors, thereby establishing a principled foundation for error broadcasting. This theoretically sound mechanism naturally leads to the experimentally observed three-factor learning rule and integrates with biologically plausible frameworks to enhance performance and plausibility. Numerical experiments demonstrate EBD’s competitive or better performance against other error-broadcast methods on benchmark datasets. Our findings establish EBD as an efficient, biologically plausible, and principled alternative for neural network training.
@@ 1 Introduction
Neural networks are dominant mathematical models for biological and artificial intelligence. A major challenge in these networks is determining how to adjust individual synaptic weights to optimize a global learning objective, which is known as the credit assignment problem. In Artificial Neural Networks (ANNs), the most common solution to this problem is the backpropagation (BP) algorithm [1].
@@
In contrast to artificial neural networks (ANNs), the mechanisms for credit assignment in biological neural networks remain poorly understood. While backpropagation is highly effective for training ANNs, it is not directly applicable to biological systems because it relies on biologically implausible assumptions. In its standard form, backpropagation propagates output errors backward through a separate pathway, reusing the same synaptic weights as in the forward pass (Figure 1a). This requirement for weight symmetry, where identical values are used for both signal transmission and error feedback, is not supported by biological evidence [2]. Although many experimentally motivated models of local synaptic plasticity have been proposed [3], a biologically feasible theory of credit assignment that integrates these mechanisms remains unresolved.
@@
To address the credit assignment problem in biological networks, researchers have proposed methods known as error broadcasting [4, 5, 6, 7, 8, 9]. These methods broadcast errors without precise backward paths or symmetric weights, eliminating the symmetry issue inherent in backpropagation. In addition to their relevance for modeling biological systems, error broadcasting offers practical benefits for hardware implementation. Recent work by [10], for example, demonstrates their potential for efficient neural network execution on hardware, offering promising directions for neuromorphic systems. However, despite encouraging progress in both theory and application [11, 12], error broadcasting still needs stronger theoretical foundations to fully validate and enhance its training effectiveness.
@@
In this context, we introduce a novel learning framework termed the Error Broadcast and Decorrelation (EBD). The fundamental principle of EBD is to establish layer-specific objectives by adjusting network weights to minimize the correlation between broadcast output errors and the activations of each layer. This approach is rigorously grounded in the theory of Minimum Mean Square Error (MMSE) estimation. A central property of an optimal MMSE estimator is that its estimation error is orthogonal to any measurable function of its input. Recognizing that each network layer computes activations which are themselves functions of the overall network input, we leverage this MMSE orthogonality principle as a principled foundation for EBD. It allows for the definition of layer-specific training losses aimed at driving these activations towards orthogonality with the broadcast error. This formulation of localized objectives at different network stages enables a more distributed mechanism for credit assignment, offering an alternative to approaches that rely solely on an output-defined loss and end-to-end error propagation.
@@
EBD directly broadcasts output errors to layers, simplifying credit assignment and enabling parallel synaptic updates. It has two key advantages for providing a framework for biologically realistic networks. First, optimizing the loss function of EBD naturally leads to the general form of experimentally observed three-factor learning rules [13, 14]. Second, by broadcasting errors directly to the layers as shown in Figure 1b, it overcomes the weight transport problem inherent in backpropagation and some more biologically plausible credit assignment approaches [15, 16].
We demonstrate EBD’s utility by applying it to both artificial and biologically realistic neural networks. Benchmark results show EBD matching/exceeding state-of-the-art error-broadcast techniques. Its successful application to a 10-layer biologically plausible network (CorInfoMax-EBD) provides initial evidence of depth scalability, offering an empirical foundation for broader investigations into scalability on more complex tasks.
@@ 1.1 Related work and contributions
Several frameworks have been proposed as alternatives to the backpropagation algorithm for modeling credit assignment in biological networks [8]. These include predictive coding [17, 15, 18], similarity matching [16, 19], time-contrastive approaches [20, 21, 22], forward-only methods [23, 24, 25], target propagation [26, 27, 28], random feedback alignment [29], and learned feedback weights [30, 31].
Another significant alternative is error-broadcast methods, where output errors are directly transmitted to network layers without relying on precise backward pathways or symmetric weights. Two important examples of this approach are weight and node perturbation algorithms [4, 32, 33, 34], in which global error signals are broadcast to all network units. These signals reflect the change in overall error caused by individual perturbations in the network’s weights or units. A more recent and prominent example of error broadcast is Direct Feedback Alignment (DFA) [6]. In DFA, the output errors are projected onto the hidden layers through fixed random weights, effectively replacing the symmetric backward weights required in traditional backpropagation. This approach first emerged as a modification to the feedback alignment approach (which replaced the symmetric weights of the backpropagation algorithm with random ones). DFA has been extended and analyzed in several studies [35, 36, 12, 37, 11], demonstrating its potential in training neural networks with less biologically implausible mechanisms. Clark et al. [9] introduced another broadcast approach for a network with vector units and nonnegative weights for which three factor learning based update rule is applied.
@@ 2.1 Problem statement
for $k=1,\ldots,L$, where $f^{(k)}$ are activation functions, $\mathbf{W}^{(k)}$ weights, and $\mathbf{b}^{(k)}$ biases. We consider input-output pairs $({\mathbf{x}},{\mathbf{y}})$ sampled from a joint distribution $P({\mathbf{x}},{\mathbf{y}})$. The performance criterion is the mean square of the output error $\bm{\epsilon}={\mathbf{h}}^{(L)}-{\mathbf{y}}$, i.e., $\mathbb{E}_{P({\mathbf{x}},{\mathbf{y}})}[\|\bm{\epsilon}\|^{2}_{2}]$.
@@ 2.2 Error Broadcast and Decorrelation loss functions
    since each hidden‐layer activation is a deterministic nonlinear function of the input ${\mathbf{x}}$, the output error of an optimal estimator should be stochastically orthogonal to those activations. Figure 1c illustrates this phenomenon by showing the evolution of the average absolute correlation between layer activations and the error signal during backpropagation training of an MLP with three hidden layers on the CIFAR-10 dataset, based on the MSE criterion. The declining correlation during MSE training reflects the MMSE estimator’s stochastic orthogonality of layer activations and output errors,
@@
    if hidden‐layer activations form a “rich enough” set of functions of ${\mathbf{x}}$ (as elaborated below), then enforcing orthogonality to them implies orthogonality to ”every” function of ${\mathbf{x}}$.
@@
$\mathbf{R}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}=\mathbb{E}[{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})\bm{\epsilon}^{T}]=\mathbf{0},\hskip 7.22743ptk=0,\ldots,L,\mbox{ with the typical choce }{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})={\mathbf{h}}^{(k)}.$ (3)
@@
Specifically, based on the stochastic orthogonality condition in Eq. (3), we propose minimizing the Frobenius norm of the cross-correlation matrices $\mathbf{R}_{{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)}),\bm{\epsilon}}$ as a replacement for the standard MSE loss. To this end, we define the estimated cross-correlation matrix between a function $g^{(k)}$ of layer activations and the output error for batch $m$ and layer $k$ as
@@
where $\lambda\in[0,1]$ is the forgetting factor used in the autoregressive estimation, $B$ is the batch size, $\hat{\mathbf{R}}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}[0]$ is the initial value hyperparameter for the correlation matrix, and
@@
is the matrix of nonlinearly transformed activations of layer $k$ for batch $m$, while
@@ 2.3 Error Broadcast and Decorrelation algorithm
${\mathbf{q}}^{(k)}[m]=\hat{\mathbf{R}}_{{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)}),\,\bm{\epsilon}}[m]\,\bm{\epsilon}[m],$
representing the projection of the output error onto the layer activations, with the cross-correlation matrix $\hat{\mathbf{R}}_{{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)}),\,\bm{\epsilon}}[m]$ as the transformation matrix. These projections are shown in Figure 1b.
@@ 2.4.1 Avoiding collapse
where $\text{where }\hskip 21.68121pt{\mathbf{H}}^{(k)}[m]$ $=[\begin{array}[]{ccc}{\mathbf{h}}^{(k)}[mB+1]&\dots&{\mathbf{h}}^{(k)}[(m+1)B]\end{array}],$ (14)
@@ 3.1 Three factor learning rule and EBD
The three-factor learning rule proposed for biological neural networks extends the traditional two-factor Hebbian rule by incorporating a modulatory signal into synaptic updates based on presynaptic and postsynaptic activity [42, 13]. Upon closer examination of the EBD update expression in Eq. (9), we observe that it conforms to the three-factor update form:
@@
[…]
@@ 3.2 CorInfoMax-EBD: CorInfoMax with three factor learning rule
$\mathcal{J}_{CI}[m]=\sum_{k=1}^{L-1}({\hat{I}_{r}^{(\varepsilon_{k})}}({\mathbf{h}}^{(k-1)},{\mathbf{h}}^{(k)})[m]+{\hat{I}_{l}^{(\varepsilon_{k})}}({\mathbf{h}}^{(k)},{\mathbf{h}}^{(k+1)})[m]),\hskip 14.45377pt\text{where,}$,
@@
Step 1: Update error projection weights: ${\mathbf{R}}_{{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})\bm{\epsilon}}[m]=\lambda_{d}{\mathbf{R}}_{{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})\bm{\epsilon}}[m-1]+\frac{1-\lambda_{d}}{B}{\mathbf{G}}^{(k)}[m]{\mathbf{E}}[m]^{T}$
Step 2: Project errors to layer $k$: ${\mathbf{Q}}^{(k)}[m]={\mathbf{R}}^{(k)}_{{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})\bm{\epsilon}}[m]{\mathbf{E}}[m]$
@@
Here, we assume layer activations ${\mathbf{H}}^{(k)}$, forward (backward) prediction errors $\overset{\rightarrow}{{\mathbf{E}}}$ ( $\overset{\leftarrow}{{\mathbf{E}}}$), output error ${\mathbf{E}}$ and lateral weight outputs ${\mathbf{Z}}$ are computed by the CorInfoMax network dynamics specified in [41] (see also Appendix F). By integrating EBD, we enable a single-phase update per input, eliminating the less biologically plausible two-phase learning mechanism required by CorInfoMax-EP. The two-phase approach of EP—comprising separate label-free and label-connected phases—is considered less plausible because biological neurons are unlikely to alternate between distinct global phases for learning. Our method not only simplifies the learning process but also aligns more closely with biological learning. Additionally, we achieve comparable or even superior performance to CorInfoMax-EP (see Section 4).

### v3

@@ Abstract
We introduce Error Broadcast and Decorrelation (EBD), a novel learning framework for neural networks that addresses credit assignment by directly broadcasting output errors to individual layers, circumventing weight transport of backpropagation. EBD is rigorously grounded in the stochastic orthogonality property of Minimum Mean Square Error estimators. This fundamental principle states that the error of an optimal estimator is orthogonal to functions of the input. Guided by this insight, EBD defines layerwise loss functions that directly penalize correlations between layer activations and output errors, thereby establishing a principled foundation for error broadcasting. This theoretically sound mechanism naturally leads to the experimentally observed three-factor learning rule and integrates with biologically plausible frameworks to enhance performance and plausibility. Numerical experiments demonstrate EBD’s competitive or better performance against other error-broadcast methods on benchmark datasets. Our findings establish EBD as an efficient, biologically plausible, and principled alternative for neural network training. The implementation is available at: https://github.com/meterdogan07/error-broadcast-decorrelation.
@@ 1 Introduction
Neural networks are dominant mathematical models for biological and artificial intelligence. A major challenge in these networks is determining how to adjust individual synaptic weights to optimize a global learning objective, known as the credit assignment problem. In Artificial Neural Networks (ANNs), the most common solution is the backpropagation (BP) algorithm [1].
@@
In contrast to ANNs, the mechanisms for credit assignment in biological neural networks remain poorly understood. While backpropagation is highly effective for training ANNs, it is not directly applicable to biological systems because it relies on biologically implausible assumptions. In its standard form, backpropagation propagates output errors backward through a separate pathway, reusing the same synaptic weights as in the forward pass (Figure 1a). This requirement for weight symmetry is not supported by biological evidence [2]. Although many experimentally motivated models of local synaptic plasticity have been proposed [3], a biologically feasible theory of credit assignment that integrates these mechanisms remains unresolved.
@@
To address the credit assignment problem in biological networks, researchers have proposed methods known as error broadcasting [4, 5, 6, 7, 8, 9]. These methods involve broadcasting the global output error directly to all layers, often through random projections or fixed pathways, without relying on precise backward paths or symmetric weights (as summarized in Section 1.1). This eliminates the weight symmetry issue inherent in backpropagation. Error broadcasting offers practical benefits for hardware implementation; recent work [10] demonstrates potential for efficient neural network execution. However, despite encouraging progress in both theory and application [11, 12], error broadcasting still needs stronger theoretical foundations to fully validate and enhance its training effectiveness.
@@
In this context, we introduce a novel learning framework termed the Error Broadcast and Decorrelation (EBD), which builds on basic error broadcasting by introducing layer-specific objectives grounded in estimation theory. The fundamental principle of EBD is to adjust network weights to minimize the correlation between broadcast output errors and the activations of each layer. This approach is rigorously grounded in Minimum Mean Square Error (MMSE) estimation, where an optimal estimator’s error is orthogonal to any measurable function of its input. We leverage this orthogonality principle for EBD, defining layer-specific training losses to drive layer activations (functions of the network input) towards orthogonality with the broadcast error. This enables a more distributed mechanism for credit assignment, alternative to approaches relying solely on an output-defined loss and end-to-end error propagation.
@@
EBD directly broadcasts output errors to layers, simplifying credit assignment and enabling parallel synaptic updates. It offers two key advantages for biologically realistic networks. First, optimizing EBD’s loss naturally leads to experimentally observed three-factor learning rules [13, 14], which extend Hebbian plasticity by incorporating a neuromodulatory signal (the third factor) modulating synaptic updates based on pre- and postsynaptic activity. Second, by broadcasting errors directly to layers as shown in Figure 1b, it overcomes the weight transport problem inherent in backpropagation and some more biologically plausible credit assignment approaches [15, 16].
We demonstrate EBD’s utility by applying it to both artificial and biologically realistic neural networks. Benchmark results show EBD matching/exceeding state-of-the-art error-broadcast techniques. Its successful application to a 10-layer biologically plausible network (CorInfoMax-EBD) provides initial evidence of depth scalability for more complex tasks.
@@ 1.1 Related work and contributions
Several frameworks have been proposed as alternatives to the backpropagation algorithm for modeling credit assignment in biological networks [8]. These include predictive coding [17, 15, 18], similarity matching [16, 19], time-contrastive approaches [20, 21, 22], forward-only methods [23, 24, 25], target propagation [26, 27, 28], random feedback alignment [29], and learned feedback weights [30, 31]. Alternative strategies also seek to establish local learning rules by optimizing statistical objectives, such as the Hilbert-Schmidt Independence Criterion (HSIC) bottleneck [32].
Another significant alternative is error-broadcast methods, where output errors are directly transmitted to network layers without relying on precise backward pathways or symmetric weights. Two important examples of this approach are weight and node perturbation algorithms [4, 33, 34, 35], in which global error signals are broadcast to all network units. These signals reflect the change in overall error caused by individual perturbations in the network’s weights or units. A more recent and prominent example of error broadcast is Direct Feedback Alignment (DFA) [6]. In DFA, the output errors are projected onto the hidden layers through fixed random weights, effectively replacing the symmetric backward weights required in traditional backpropagation. The core challenge of weight transport has been tackled by several other methods, many of which also rely on fixed random signals or avoid feedback entirely [36, 37]. Encouragingly, a number of these biologically-plausible frameworks have demonstrated the ability to scale effectively to large datasets, underscoring their potential as viable training mechanisms [38]. This approach first emerged as a modification to the feedback alignment approach (which replaced the symmetric weights of the backpropagation algorithm with random ones). DFA has been extended and analyzed in several studies [39, 40, 12, 41, 11], demonstrating its potential in training neural networks with less biologically implausible mechanisms. Clark et al. [9] introduced another broadcast approach for a network with vector units and nonnegative weights for which three factor learning based update rule is applied.
@@ 2.1 Problem statement
where $k\in\{1,\ldots L\}$ is the layer index, $f^{(k)}$ are activation functions, $\mathbf{W}^{(k)}$ weights, ${\mathbf{u}}^{(k)}$ preactivations and $\mathbf{b}^{(k)}$ biases. We consider input-output pairs $({\mathbf{x}},{\mathbf{y}})$ sampled from a joint distribution $P({\mathbf{x}},{\mathbf{y}})$. The performance criterion is the mean square of the output error $\bm{\epsilon}={\mathbf{h}}^{(L)}-{\mathbf{y}}$, i.e., $\mathbb{E}_{P({\mathbf{x}},{\mathbf{y}})}[\|\bm{\epsilon}\|^{2}_{2}]$.
@@ 2.2 Error Broadcast and Decorrelation loss functions
    Since each hidden-layer activation is a nonlinear function of input $\mathbf{x}$}, the output error of an optimal estimator should be stochastically orthogonal to those activations. Figure 1c illustrates this phenomenon by showing the evolution of the average absolute correlation between layer activations and the error signal during backpropagation training of an MLP with three hidden layers on the CIFAR-10 dataset, based on the MSE criterion. Similar correlation declining trends are also observed across different datasets and architectures (see Appendix J). The declining correlation during MSE training reflects the MMSE estimator’s stochastic orthogonality of layer activations and output errors.,
@@
    if hidden‐layer activations form a “rich enough” set of functions of $\mathbf{x}$} (as elaborated below), then enforcing error orthogonality to them implies orthogonality to ”every” function of $\mathbf{x}$}.
@@
$\mathbf{R}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}=\mathbb{E}[{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})\bm{\epsilon}^{T}]=\mathbf{0},\hskip 7.22743ptk=0,\ldots,L,\mbox{ with the typical choice }{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})={\mathbf{h}}^{(k)}.$ (3)
@@
Specifically, based on the stochastic orthogonality condition in Eq. (3), we propose minimizing the Frobenius norm of the cross-correlation matrices $\mathbf{R}^{(k)}_{{\mathbf{g}}\bm{\epsilon}}$ as a replacement for the standard MSE loss. To this end, we define the estimated cross-correlation matrix between a function $g^{(k)}$ of layer activations and the output error for batch $m$ and layer $k$ as
@@
where $m$ is the batch index, $\lambda\in[0,1]$ is the forgetting factor used in the autoregressive estimation, $B$ is the batch size, $\hat{\mathbf{R}}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}[0]$ is the initial value hyperparameter for the correlation matrix, and
@@
is the matrix of nonlinearly transformed activations of layer $k$ for batch $m$. In the above equation, $mB+l$ refers to absolute (sequence) index for the $l^{th}$ member of batch-$m$. Furthermore,
@@ 2.3 Error Broadcast and Decorrelation algorithm
${\mathbf{q}}^{(k)}[m]=\hat{\mathbf{R}}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}[m]\,\bm{\epsilon}[m],$
representing the projection of the output error onto the layer activations, with the cross-correlation matrix $\hat{\mathbf{R}}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}[m]$ as the transformation matrix. These projections are shown in Figure 1b.
For the special case of batchsize, $B=1$, the weight update in (9) simplifies to
$[\Delta{\bm{W}}^{(k)}_{1}[m]]_{ij}=\zeta{g^{\prime}}^{(k)}_{i}(h_{i}^{(k)}[m]){f^{\prime}}^{(k)}(u_{i}^{(k)}[m])q_{i}^{(k)}[m]h_{j}^{(k-1)}[m].$ (10)
@@ 2.4.1 Avoiding collapse
$\text{where }\hskip 21.68121pt{\mathbf{H}}^{(k)}[m]$ $=[\begin{array}[]{ccc}{\mathbf{h}}^{(k)}[mB+1]&\dots&{\mathbf{h}}^{(k)}[(m+1)B]\end{array}],$ (16)
@@ 3 EBD for biologically realistic networks
Figure 2: Error–broadcast learning as a three‑factor synaptic update. The presynaptic firing rate $h^{(k-1)}_{j}$ (green, left) projects onto the postsynaptic neuron $h^{(k)}_{i}$ (blue, centre) through the synapse $W^{(k)}_{ij}$. A layer‑specific broadcast of the output error $e\!\to\!q^{(k)}_{i}$ (yellow, right) provides the modulatory third factor that gates plasticity. Together, presynaptic activity, postsynaptic non‑linear derivatives $g^{\prime(k)}_{i}f^{\prime(k)}$ (blue rectangle), and the modulatory signal form the product that drives the EBD weight change $\Delta W^{(k)}_{ij}$ displayed underneath the circuit.
@@ 3.1 Three factor learning rule and EBD
The three-factor learning rule for biological neural networks extends the traditional two-factor Hebbian rule by incorporating a modulatory signal into synaptic updates based on presynaptic and postsynaptic activity [46, 13]. While backpropagation can be expressed similarly, it is not typically considered a three-factor rule in neuroscience, as its ’third factor’ is a locally tailored signal specific to each neuron requiring a biologically implausible dual network with symmetric weights, unlike global neuromodulatory signals [13]. In contrast, EBD update for batchsize $B=1$ in (10) naturally matches the three-factor structure:
@@
where $q_{i}^{(k)}$ is the projected global error. Thus, EBD supports various three-factor rules depending on nonlinearity $g^{(k)}$. For example, $g_{i}^{(k)}(h_{i}^{(k)})={h_{i}^{(k)}}^{2}$ yields the error-modulated Hebbian update [47, 11]: $\Delta W_{ij}^{(k)}\propto{h_{i}^{(k)}{f^{\prime}}^{(k)}(u_{i}^{(k)})}\hskip 1.49994pt{q_{i}^{(k)}}\hskip 1.49994pt{h_{j}^{(k-1)}}$. By enabling diverse three-factor updates via different nonlinear functions, EBD holds potential for modeling biologically consistent neural learning processes.
Figure 2 breaks the EBD weight update (10) into its three interacting factors. The presynaptic activity $h^{(k-1)}_{j}$ from the sending unit; the postsynaptic term is $g_{i}^{\prime(k)}f^{\prime(k)}$ computed from the receiving unit’s own activation; and the modulatory broadcast error $q_{i}^{(k)}$, derived from the network’s output error $\bm{\epsilon}$. Multiplying these three quantities produces the weight change $\Delta W^{(k)}_{ij}$ shown beneath the diagram, revealing that EBD naturally realises the classical three‑factor learning rule in neural networks.
@@ 3.2 CorInfoMax-EBD: CorInfoMax with three factor learning rule
$\mathcal{J}_{CI}[m]=\sum_{k=1}^{L-1}({\hat{I}_{r}^{(\varepsilon_{k})}}({\mathbf{h}}^{(k-1)},{\mathbf{h}}^{(k)})[m]+{\hat{I}_{l}^{(\varepsilon_{k})}}({\mathbf{h}}^{(k)},{\mathbf{h}}^{(k+1)})[m]),\hskip 14.45377pt\text{where,}$
@@
Step 1: Update error projection weights: $\hat{\mathbf{R}}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}[m]=\lambda_{d}\hat{\mathbf{R}}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}[m-1]+\frac{1-\lambda_{d}}{B}{\mathbf{G}}^{(k)}[m]{\mathbf{E}}[m]^{T}$
Step 2: Project errors to layer $k$: ${\mathbf{Q}}^{(k)}[m]=\hat{\mathbf{R}}_{{\mathbf{g}}\bm{\epsilon}}^{(k)}[m]{\mathbf{E}}[m]$
@@
Here, we assume layer activations ${\mathbf{H}}^{(k)}$, forward (backward) prediction errors $\overset{\rightarrow}{{\mathbf{E}}}$ ($\overset{\leftarrow}{{\mathbf{E}}}$), output error $\mathbf{E}$}, and lateral weight outputs $\mathbf{Z}$} are computed via CorInfoMax network dynamics in [45] (see Appendix F). Integrating EBD enables single-phase updates per input, eliminating the less biologically plausible two-phase mechanism required by CorInfoMax-EP. EP’s two-phase approach—separate label-free and label-connected phases—is implausible, as biological neurons unlikely alternate between distinct global phases for learning. Our method simplifies the process, aligns more closely with biological learning, and achieves comparable or superior performance to CorInfoMax-EP (see Section 4).